OpenStudy (anonymous):

Radicals!!! Approximate (to the nearest tenth of an inch) the length of the missing side of the following right triangles. a^2+4^2=17^2

OpenStudy (anonymous):

\[a^2+4^2=17^2 \implies a^2=17^2-4^2 \implies a^2=285 \implies a=\sqrt{285}.\] Use a calculator to find that \(a = 16.5\).